# LTGC: Long-tail Recognition via Leveraging LLMs-driven Generated Content

### 09 Mar 2024 v1

@@ 5 Conclusion
Table A3: Effectiveness of the Self-Checking module. NC: Number-checking. RC: Repetition-checking.
@@
Effectiveness of the self-checking module. The Self-Checking module consists of two elements: a number check for images and a repetition check. We separately investigated these two checks, and the results are shown in Tab. A3. When only performing the quality check, LLMs generate highly repetitive descriptions, leading to a decrease in textual diversity and, consequently, a decline in image diversity. When only performing the repetitiveness check, the LLMs are posed the two questions mentioned in Section 3.1 only once. This results in a limited number of generated samples, thereby leading to limited performance gains. As shown in Tab. A3, our proposed method incorporates both checks and consistently outperforms all variants. This attests to (1) the effectiveness of the repetitiveness design, enabling LLMs to generate comprehensive text descriptions, and (2) the efficacy of the self-checking design, allowing LLMs to increase the number of samples and simulate a more balanced set of classes.
@@
The effectiveness of different maximum image numbers for LLM’s self-checking module. In our framework, the maximum number of generated and original images for each class $y$, denoted as $K_{y}$. For LLM’s self-checking module, the $K_{y}$ is set to 100 for iNaturalist 2018, 300 for ImageNet-LT, and 800 for Place-LT, respectively. We explore alternative caps for the maximum image number and show the findings in Fig. A1. The results indicate that for each dataset when the cap is set below 100, 300, and 800, respectively, there’s a noticeable improvement in our framework’s performance as the limit on generative images increases. This might be because, by setting the maximum number of images to be a larger number, LLM can better cover the diverse distinctive features and backgrounds (scenes). Moreover, it’s observed that increasing the maximum image number beyond 100, 300, and 800 does not lead to further improvements in performance. However, surpassing these maximum numbers of 100, 300, and 800 doesn’t yield further performance gains. Therefore, balancing performance and efficiency, we determined these respective limits as the optimal settings for our framework.
Figure A1: The effectiveness of different maximum image numbers for LLM’s self-checking module.
Impact of the maximum number of cycle times for cyclic-assessing module. In our framework’s cyclic-assessing module, we initially cap the cycle count at three. To explore the impact of varying this limit, we experimented with different maximum cycle thresholds and presented the outcomes in Table A5. The results indicate an enhancement in our framework’s performance as we expand the maximum cycle count for cyclic cross-assessing. This improvement is likely due to the more thorough refinement of generated images achieved with a higher cycle limit. However, the gains in performance diminish when the cycle count exceeds three, leading us to establish three as the optimal maximum for cyclic-assessing module.
Table A5: Evaluation on the maximum number of cycle times for cyclic-assessing.
@@
Check the reasonability of the generated descriptions. In our framework, we use ChatGPT to generate text descriptions for the evaluated classes, and we further design a cyclic cross-assessing module in which we guide ChatGPT to modify its generated descriptions. Here, we perform a check w.r.t. the reasonability of the generated descriptions before and after passing into the cross-assessing module. Specifically, we find that, before passing into the cross-assessing module, 4% of descriptions are checked to be unreasonable, but 0.2% of the descriptions output by the module is checked to be unreasonable. This shows that ChatGPT has small probability of generating unreasonable descriptions, while our cyclic-assessing module can further mitigate this problem. The above check is done by inviting 3 volunteers and passing the same 1000 descriptions to them. The 3 volunteers first make decisions independently and then discuss disagreed decisions.
Figure A2: Visualization about the cyclic cross-assessing module. The images in (a) do not align most closely with their intended classes until the third cycle, whereas the images in (b) achieve their closest alignment with the intended classes in the second cycle. The classes of Tamias Amoenus and Lepus Townsendii are derived from the iNaturalist dataset, while Swimming Pool and Triumphal Arch are from the Place-LT and ImageNet-LT datasets, respectively.
@@
Visualisation of the cyclic-assessing module. In this work, we propose a cyclic cross-assessing module in our framework. This module aims to enhance the quality of generated images, particularly those that are not highly accurate. Fig A2 illustrates the refinement process carried out by our cyclic-assessing module. As depicted, this module effectively improves images that initially do not align well with their intended classes, ensuring they more accurately represent these classes. This highlights the efficiency and utility of our cross-assessing module in image refinement.

### 13 Mar 2024 v3

@@ 5 Conclusion
Table A3: Effectiveness of the Self-reflection module. NC: Number-checking. RC: Repetition-checking.
@@
Effectiveness of the Self-reflection module. The self-reflection module consists of two elements: a number check for images and a repetition check. We separately investigated these two checks, and the results are shown in Tab. A3. When only performing the quality check, LLMs generate highly repetitive descriptions, leading to a decrease in textual diversity and, consequently, a decline in image diversity. When only performing the repetitiveness check, the LLMs are posed the two questions mentioned in Section 3.1 only once. This results in a limited number of generated samples, thereby leading to limited performance gains. As shown in Tab. A3, our proposed method incorporates both checks and consistently outperforms all variants. This attests to (1) the effectiveness of the repetitiveness design, enabling LLMs to generate comprehensive text descriptions, and (2) the efficacy of the self-reflection design, allowing LLMs to increase the number of samples and simulate a more balanced set of classes.
@@
The effectiveness of different maximum image numbers for LLM’s self-reflection module. In our framework, the maximum number of generated and original images for each class $y$, denoted as $K_{y}$. For LLM’s self-reflection module, the $K_{y}$ is set to 100 for iNaturalist 2018, 300 for ImageNet-LT, and 800 for Place-LT, respectively. We explore alternative caps for the maximum image number and show the findings in Fig. A1. The results indicate that for each dataset when the cap is set below 100, 300, and 800, respectively, there’s a noticeable improvement in our framework’s performance as the limit on generative images increases. This might be because, by setting the maximum number of images to be a larger number, LLM can better cover the diverse distinctive features and backgrounds (scenes). Moreover, it’s observed that increasing the maximum image number beyond 100, 300, and 800 does not lead to further improvements in performance. However, surpassing these maximum numbers of 100, 300, and 800 doesn’t yield further performance gains. Therefore, balancing performance and efficiency, we determined these respective limits as the optimal settings for our framework.
Figure A1: The effectiveness of different maximum image numbers for LLM’s self-reflection module.
Impact of the maximum number of cycle times for iterative evaluation module. In our framework’s iterative evaluation module, we initially cap the cycle count at three. To explore the impact of varying this limit, we experimented with different maximum cycle thresholds and presented the outcomes in Table A5. The results indicate an enhancement in our framework’s performance as we expand the maximum cycle count for iterative evaluation. This improvement is likely due to the more thorough refinement of generated images achieved with a higher cycle limit. However, the gains in performance diminish when the cycle count exceeds three, leading us to establish three as the optimal maximum for the iterative evaluation module.
Table A5: Evaluation on the maximum number of cycle times for iterative evaluation.
@@
Check the reasonability of the generated descriptions. In our framework, we use ChatGPT to generate text descriptions for the evaluated classes, and we further design a iterative evaluation module in which we guide ChatGPT to modify its generated descriptions. Here, we perform a check w.r.t. the reasonability of the generated descriptions before and after passing into the iterative evaluation module. Specifically, we find that, before passing into the iterative evaluation module, 4% of descriptions are checked to be unreasonable, but 0.2% of the descriptions output by the module is checked to be unreasonable. This shows that ChatGPT has small probability of generating unreasonable descriptions, while our iterative evaluation module can further mitigate this problem. The above check is done by inviting 3 volunteers and passing the same 1000 descriptions to them. The 3 volunteers first make decisions independently and then discuss disagreed decisions.
Figure A2: Visualization about the iterative evaluation module. The images in (a) do not align most closely with their intended classes until the third cycle, whereas the images in (b) achieve their closest alignment with the intended classes in the second cycle. The classes of Tamias Amoenus and Lepus Townsendii are derived from the iNaturalist dataset, while Swimming Pool and Triumphal Arch are from the Place-LT and ImageNet-LT datasets, respectively.
@@
Visualisation of the iterative evaluation module. In this work, we propose an iterative evaluation module for our framework. This module aims to enhance the quality of generated images, particularly those that are not highly accurate. Fig A2 illustrates the refinement process carried out by our iterative evaluation module. As depicted, this module effectively improves images that initially do not align well with their intended classes, ensuring they more accurately represent these classes. This highlights the efficiency and utility of our iterative evaluation module in image refinement.